OpenStudy (anonymous):

Help? Which is the inverse of y=1/2x+3? A y=2x−3 B y=2x−6 C y=−3x + 2 D y=−1/2 x − 3

OpenStudy (luigi0210):

switch the x and y and solve for y

OpenStudy (anonymous):

so D?

OpenStudy (luigi0210):

Not exactly.. y=1/2x+3 To find the inverse switch the variables: x=1/2y+3

OpenStudy (luigi0210):

So solve for y: x-3=1/2x multiply by 2 2x-6=y

OpenStudy (johnweldon1993):

And now solve again for y. so first isolate 1/2y on the right side and then COMPLETELY isolate y to solve for y again

OpenStudy (anonymous):

so, B

OpenStudy (johnweldon1993):

correct!

OpenStudy (luigi0210):

Yup, yup.

OpenStudy (anonymous):

thanks!

OpenStudy (luigi0210):

Any time :D